OpenStudy (anonymous):

(a) Calculate the torque produced by a 40 N perpendicular force at the end of a 0.1 m long wrench.

OpenStudy (anonymous):

4

OpenStudy (anonymous):

u just multiply the numbers together

OpenStudy (anonymous):

omg so i am Will-i-am except that someone got my account deleted plz help

OpenStudy (anonymous):

torque (t) is defined as the force (N) applied multiplied by the level arm length (L) at a perpendicular angle. In this case, you have the equation t = F * L sin (90 deg) = .1m * 40N = 4 Nm

OpenStudy (anonymous):

since its a perpendicular force t =f*radius =40*0.1=4